alright and the last one for tonight !
Jim can fill a pool carrying buckets of water in 30mins. Sue can do the same job in 45mins. Tony can do the same job in 1 1/2 hours. How quickly can all three fill the pool together?
a.12 minutes
b.15 minutes
c.23minutes
d.28 minutes
e.21 minutes
thank you and good night!

Answer by ptaylor(2198)


OK, we'll deal in minutes
Let x=amount of time it takes all three working together to fill the pool
Jim fills at the rate of 1/30 pool per min
Sue fills at the rate of 1/45 pool per min
Tony fills at the rate of 1/90 pool per min
Together they fill at the rate of 1/30+1/45+1/90=
3/90 + 2/90 + 1/90=6/90 pool per min
So, our equation to solve is
(6/90)x=1(1 pool that is) multiply each side by 90
6x=90 divide each side by 6
x=15 min-------------------time it takes all three working together

Hope this helps---ptaylor